If a classroom has a length that is 5 feet less than two times its width and its perimeter is 80 feet,find width and length of the classroom?
Answer by vheroli(126)

Formulas:
P= 2L + 2W
L= 2x - 5
W= x

Solution:
80 = [ 2(2x-5) ] + 2(x) :Simplify
80 = (4x-10) + 2x : Simplify
80 = 6x - 10 : Simplify
80 + 10 = 6x
90 = 6x : Solve for x
90/6 = x
15 = x

Answers:
L= 2x - 5
L= 2 (15) - 5
L= 30-5
L= 25 ft

W= x
W= 15 ft